# Uncertainty-Aware Planning for Heterogeneous Robot Teams using Dynamic Topological Graphs and Mixed-Integer Programming

### v4

@@ Abstract
Multi-robot planning and coordination in uncertain environments is a fundamental computational challenge, since the belief space increases exponentially with the number of robots. In this paper, we address the problem of planning in uncertain environments with a heterogeneous robot team of fast scout vehicles for information gathering and more risk-averse carrier robots from which the scouts vehicles are deployed. To overcome the computational challenges, we represent the environment and operational scenario using a topological graph, where the parameters of the edge weight distributions vary with the state of the robot team on the graph, and we formulate a computationally efficient mixed-integer program which removes the dependence on the number of robots from its decision space. Our formulation results in the capability to generate optimal multi-robot, long-horizon plans in seconds that could otherwise be computationally intractable. Ultimately our approach enables real-time re-planning, since the computation time is significantly faster than the time to execute one step. We evaluate our approach in a scenario where the robot team must traverse an environment while minimizing detection by observers in positions that are uncertain to the robot team. We demonstrate that our approach is computationally tractable, can improve performance in the presence of imperfect information, and can be adjusted for different risk profiles.
@@ III PROBLEM STATEMENT
We consider the problem of planning and coordination for a heterogeneous robot team by generating a graph $G=(V,E,w_{e})$, with nodes $v\in V$ and edges $e\in E$, based on the critical features of the planning problem. The edge cost $w_{e}:E\times X\times T\rightarrow\mathbb{R}_{>0}$ is a bounded random variable that depends on the underlying environmental uncertainty $u_{e}\in[u_{e}^{\llcorner},u_{e}^{\ulcorner}]$, the discrete robot team state on the graph $x\in X$, and time $t\in\mathbb{R}_{>0}$. We consider $n_{CL}$ robot classes and $n_{E}$ total edges such that $X=X^{1}\times X^{2}\times\dots\times X^{n_{CL}}$, where $X^{i}=\{[p_{e_{1}}^{i},p_{e_{2}}^{i},\dots,p_{e_{n_{E}}}^{i}]\in\mathbb{Z}_{>0}^{n_{E}}\}$ and $p_{e_{j}}^{i}$ is the number of robots of class $i$ on edge $e_{j}$. We assume a level of abstraction where multiple robots may simultaneously traverse an edge. We also allow robots to uncover the true edge cost by inspecting an edge. Following an inspection, we assume that the edge cost uncertainty returns to its original value after a finite time.
We furthermore let each robot class $i$ be constrained to graph $G^{i}=(V^{i},E^{i},w_{e}^{i})$, where $V^{i}\subseteq V$, $E^{i}\subseteq E$, and $w_{e}^{i}(e^{i},x,t^{i})=\zeta^{i}w_{e}(e^{i},x,t^{i})$ for $e^{i}\in E^{i}$, $x\in X$, and $t^{i}\in[1,n_{\tau}^{i}]$. We can now capture heterogeneous mobility, cost, speed, and sensing characteristics by defining for each robot class a graph $G^{i}$, an edge cost scale factor $\zeta^{i}$, and traversal time factor $n_{\tau}^{i}$. We also permit a set of costs and constraints that govern inter-class interactions.
@@ III-A Reconnaissance Test Scenario
We restrict our robot team to two classes: carrier robots and scout robots. Scout robots move more quickly than carrier robots, incur less cost, and possess sensors for inspecting edges to reduce uncertainty bounds to zero. Carrier robots can carry at most one scout robot. Scouts can deploy when the carrier robots are at nodes and return within one carrier robot time increment to any empty carrier robot. When a scout vehicle is deployed from a carrier robot, it incurs a specified deployment cost. In this paper, the edge cost scale factor for carrier robots and scouts are $1$ and $\zeta$, respectively, and the traversal time factors are $1$ and $n_{\tau}$, respectively. We also assume that all robots operate on graphs with identical sets of vertices and edges, i.e., $E^{i}=E,\quad V^{i}=V,\quad\forall i$.
@@
IV MIXED INTEGER PROGRAMMING APPROACH
@@
$C_{U_{A_{e,t}}}$ Cont 0 $\infty$ Carrier cost of uncertainty for $e$
$C_{U_{K_{e^{\prime},t^{\prime}}}}$ Cont 0 $\infty$ Scout cost of uncertainty for $e^{\prime}$
@@ IV-A Mathematical Preliminaries
Proposition 1: If $\alpha\in\{0,1\}$ is a binary variable, $b\in\mathbb{R}$, and $h(b)$ is a linear function with bounds $h^{\llcorner}\leq h(b)\leq h^{\ulcorner}$, then minimizing $\alpha h(b)$ is equivalent to $\min y$ for slack variable $y\in\mathbb{R}$ with the following constraints.
@@
Corollary 1: When minimizing, if all other constraints are independent of $y$ and the coefficient of $y$ in the cost function is nonnegative, then only the left side constraints are structural, and thus necessary for optimization.
@@
Special Case 1: For a constant $a\geq 0$, $h(b)=1-b$, $b\in[0,1]$, and thus $h^{\llcorner}=0$ and $h^{\ulcorner}=1$, the minimization of $a\alpha(1-b)$ can be equivalently expressed as follows.
@@
Proposition 2: For $b\in[0,b^{\ulcorner}]$, an indicator variable $\alpha\in\{0,1\}$ can be expressed as follows, for any constant $a\geq 0$.
@@ IV-B1 Cost of Traversing and Uncertainty
In our test case, our aim is to be risk-averse, so we can assume $u_{e}^{\llcorner}<u_{e}^{\ulcorner}$ and $\beta<0.5$ to ensure $\hat{u}_{e}\geq 0$ for all edges $e$. We can then apply Special Case 1 from Sec. IV-A, by using the cost terms $C_{U_{A_{e,t}}}$ and $C_{U_{K_{e,t}}}$ as slack variables with lower bounds of 00. We add the following constraints.
$C_{U_{A_{e,t}}}$ $\geq\hat{u}_{e}(\phi_{e,t}-z_{e,t})$ (17)
$C_{U_{K_{e,t}}}$ $\geq\zeta\hat{u}_{e}(\theta_{e,t}-z_{e,t})$ (18)
Following Special Case 1, $C_{U_{A_{e,t}}}$ and $C_{U_{K_{e,t}}}$ will be strict to the minimum possible values given the constraints, and thus can replace (15) and (16). Overall, this constraint formulation allows the cost of uncertainty to remain linear with linear constraints, which significantly improves solve time.
@@ IV-B2 Cost of Launch
To represent the risk associated with deploying a scout vehicle (e.g., an aerial vehicle may be noticeable and noisy), we introduce a cost for launching the vehicle, $C_{F_{v,t}}$. The launch cost for a particular vertex, $\eta_{v}$, is the average cost of the surrounding edges multiplied by a scaling factor. We express the cost for all $t$ and $v$.
@@ IV-C1 Time Tracking Variables
We define a constraint for the time tracking variables $\psi_{t}$. A robot would not pause on an edge due to the edge’s weight, thus being on an edge implies movement. In the overall cost function, $\psi_{t}$ only contributes to increasing cost, so we constrain $\psi_{t}$ using Proposition 2 from Sec. IV-A, such that $\psi_{t}=1$ when any number of robots move on an edge and otherwise $\psi_{t}=0$.
@@ IV-C7 Edge Inspection Variables
We then switch coordinates in (29) to be relative to the time $t$ and use the binary $\delta_{e,t}$ to sum components corresponding to completed inspections. In the following constraint, for all $e$ and $t$, the fractional component controls the impact of each time step’s information on the overall sum, with increasingly smaller values for less recent inspections.
@@
Both $z_{e,t}\in[0,1]$ and $\delta_{e,t}\in\{0,1\}$ only contribute to the overall cost function (8) by reducing cost so the optimizer will seek to maximize these values. Thus, equality will be achieved in both (28) and (30) up to their maximum bounds.
@@ IV-D Uncertainty-Aware Multi-Robot Optimization Problem
Scout and Uncertainty Constraints $C_{U_{A_{e,t}}}\geq\hat{u}_{e}(\phi_{e,t}-z_{e,t}),$ $\forall e,t$ (17)
$C_{U_{K_{e,t}}}\geq\hat{u}_{e}(\theta_{e,t}-z_{e,t}),$ $\forall e,t$ (18)
@@ V-A Ablation Study
Figure 3: Ablation study illustrating the advantages of considering uncertainty, scouts for the reduction of uncertainty, and certainty decay for evolving conditions. Each plot shows the final paths of the robot units through the graph. Each unit is composed of a robot carrier and scout. The first case is the baseline from [2] and each subsequent case adds one of the components from the algorithm proposed in this paper to show the resulting effect to the final team routes through the graph. In (c) and (d), the darkness of the edges reflect the number of time steps the scouts explored that edge.
@@ VI CONCLUSION
In this paper, we explored multi-robot planning on dynamic topological graphs using mixed-integer programming for two challenging cases: heterogeneous teams and planning under uncertainty. We use the uncertainty in our problem as a motivation for a heterogeneous team, introducing scout robots that can investigate the environment to collect data and reduce the uncertainty of future team actions. Our approach results in a MILP problem that can be solved rapidly with a receding horizon in real-world scenarios. We tested this approach in example scenarios and demonstrate the ability to successfully generate risk-aware plans for multi-robot teams.

### v5

@@ Abstract
Multi-robot planning and coordination in uncertain environments is a fundamental computational challenge, since the belief space increases exponentially with the number of robots. In this paper, we address the problem of planning in uncertain environments with a heterogeneous robot team of fast scout vehicles for information gathering and more risk-averse carrier robots from which the scouts vehicles are deployed. To overcome the computational challenges, we represent the environment and operational scenario using a topological graph, where the parameters of the edge weight distributions vary with the state of the robot team on the graph, and we formulate a computationally efficient mixed-integer program which removes the dependence on the number of robots from its decision space. Our formulation results in the capability to generate optimal multi-robot, long-horizon plans in seconds that could otherwise be computationally intractable. Ultimately our approach enables real-time re-planning, since the computation time is significantly faster than the time to execute one step. We evaluate our algorithm in a scenario where the robot team must traverse an environment while minimizing detection by observers in positions that are uncertain to the robot team. We demonstrate that our method is computationally tractable, can improve performance in the presence of imperfect information, and can be adjusted for different risk profiles.
@@ III PROBLEM STATEMENT
We consider the problem of planning and coordination for a heterogeneous robot team by generating a graph $G=(V,E,W_{E})$, with nodes $v\in V$ and edges $e\in E$, based on the critical features of the planning problem. The edge costs $w_{e}\in W_{E}:E\times X\times T\rightarrow\mathbb{R}_{>0}$ are bounded random variables that depend on the underlying environmental uncertainty $u_{e}\in[u_{e}^{\llcorner},u_{e}^{\ulcorner}]$, the discrete robot team state on the graph $x\in X$, and time $t\in\mathbb{R}_{>0}$. We consider $n_{CL}$ robot classes and $n_{E}$ total edges such that $X=X^{1}\times X^{2}\times\dots\times X^{n_{CL}}$, where $X^{i}=\{[p_{e_{1}}^{i},p_{e_{2}}^{i},\dots,p_{e_{n_{E}}}^{i}]\in\mathbb{Z}_{>0}^{n_{E}}\}$ and $p_{e_{j}}^{i}$ is the number of robots of class $i$ on edge $e_{j}$. We assume a level of abstraction where multiple robots may simultaneously traverse an edge. We also allow robots to uncover the true edge cost by inspecting an edge. Following an inspection, we assume that the edge cost uncertainty returns to its original value after a finite time.
We furthermore let each robot class $i$ be constrained to graph $G^{i}=(V^{i},E^{i},W_{E}^{i})$, where $V^{i}\subseteq V$, $E^{i}\subseteq E$, and $w_{e}^{i}\in W_{E}^{i}$ such that $w_{e}^{i}(e^{i},x,t^{i})=\zeta^{i}w_{e}(e^{i},x,t^{i})$ for $e^{i}\in E^{i}$, $x\in X$, and $t^{i}\in[1,n_{\tau}^{i}]$. We can now capture heterogeneous mobility, cost, speed, and sensing characteristics by defining for each robot class a graph $G^{i}$, an edge cost scale factor $\zeta^{i}$, and traversal time factor $n_{\tau}^{i}$. We also permit a set of costs and constraints that govern inter-class interactions.
@@ III-A Reconnaissance Test Scenario
We restrict our robot team to two classes: carrier robots and scout robots. Scout robots move more quickly than carrier robots, incur less cost, and possess sensors for inspecting edges to reduce uncertainty bounds to zero. Carrier robots can carry at most one scout robot. Scouts can deploy when the carrier robots are at nodes and return within one carrier robot time increment to any empty carrier robot. When a scout vehicle is deployed from a carrier robot, it incurs a specified deployment cost. In this paper, the edge cost scale factor for carrier robots and scouts are $1$ and $\zeta$, respectively, and the traversal time factors are $1$ and $n_{\tau}$, respectively. We also assume that all robots operate on graphs with identical sets of vertices and edges (i.e., $E^{i}=E,\quad V^{i}=V,\quad\forall i$).
@@
IV MIXED-INTEGER PROGRAMMING APPROACH
@@
$\breve{C}_{U_{A_{e,t}}}$ Cont 0 $\infty$ Carrier cost of uncertainty for $e$
$\breve{C}_{U_{K_{e^{\prime},t^{\prime}}}}$ Cont 0 $\infty$ Scout cost of uncertainty for $e^{\prime}$
@@ IV-A Mathematical Preliminaries
Proposition 1.
If $\alpha\in\{0,1\}$ is a binary variable, $b\in\mathbb{R}$, and $h(b)$ is a linear function with bounds $h^{\llcorner}\leq h(b)\leq h^{\ulcorner}$, then minimizing $\alpha h(b)$ is equivalent to $\min y$ for slack variable $y\in\mathbb{R}$ with the following constraints.
@@
Corollary 1.
When minimizing, if all other constraints are independent of $y$ and the coefficient of $y$ in the cost function is nonnegative, then only the left side constraints are structural, and thus necessary for optimization.
@@
Special Case 1.
For a constant $a\geq 0$, $h(b)=1-b$, $b\in[0,1]$, and thus $h^{\llcorner}=0$ and $h^{\ulcorner}=1$, the minimization of $a\alpha(1-b)$ can be equivalently expressed as follows.
@@
Proposition 2.
For $b\in[0,b^{\ulcorner}]$, an indicator variable $\alpha\in\{0,1\}$ can be expressed as follows, for any constant $a\geq 0$.
@@ IV-B1 Cost of Traversing and Uncertainty
In our test case, our aim is to be risk-averse, so we can assume $u_{e}^{\llcorner}<u_{e}^{\ulcorner}$ and $\beta<0.5$ to ensure $\hat{u}_{e}\geq 0$ for all edges $e$. We can then apply Special Case 1, by using the cost terms $\breve{C}_{U_{A_{e,t}}}$ and $\breve{C}_{U_{K_{e,t}}}$ as slack variables with lower bounds of 00. We add the following constraints.
$\breve{C}_{U_{A_{e,t}}}$ $\geq\hat{u}_{e}(\phi_{e,t}-z_{e,t})$ (17)
$\breve{C}_{U_{K_{e,t}}}$ $\geq\zeta\hat{u}_{e}(\theta_{e,t}-z_{e,t})$ (18)
Following Special Case 1, $\breve{C}_{U_{A_{e,t}}}$ and $\breve{C}_{U_{K_{e,t}}}$ will be strict to the minimum possible values given the constraints, and thus can replace (15) and (16). Overall, this constraint formulation allows the cost of uncertainty to remain linear with linear constraints, which significantly improves solve time.
@@ IV-B2 Cost of Launch
To represent the risk associated with deploying a scout vehicle (e.g., an aerial vehicle may be noticeable and noisy), we introduce a cost for launching the vehicle, $C_{F_{v,t}}$. The launch cost for a particular node, $\eta_{v}$, is the average cost of the surrounding edges multiplied by a scaling factor. We express this cost for all $t$ and $v$.
@@ IV-C1 Time Tracking Variables
We define a constraint for the time tracking variables $\psi_{t}$. A robot would not pause on an edge due to the edge’s weight, thus being on an edge implies movement. In the overall cost function, $\psi_{t}$ only contributes to increasing cost, so we constrain $\psi_{t}$ using Proposition 2, such that $\psi_{t}=1$ when any number of robots move on an edge and otherwise $\psi_{t}=0$.
@@ IV-C7 Edge Inspection Variables
We then switch coordinates of (29) to be relative to time $t$ and use the binary $\delta_{e,t}$ to sum components corresponding to completed inspections. In the following constraint, for all $e$ and $t$, the fractional component controls the impact of each time step’s information on the overall sum, with increasingly smaller values for less recent inspections.
@@
Both $z_{e,t}\in[0,1]$ and $\delta_{e,t}\in\{0,1\}$ only contribute to the overall cost function (8) by reducing cost, so the optimizer will seek to maximize these values. Thus, equality will be achieved in both (28) and (30) up to their maximum bounds.
@@ IV-D Uncertainty-Aware Multi-Robot Optimization Problem
Scout and Uncertainty Constraints $\breve{C}_{U_{A_{e,t}}}\geq\hat{u}_{e}(\phi_{e,t}-z_{e,t}),$ $\forall e,t$ (17)
$\breve{C}_{U_{K_{e,t}}}\geq\hat{u}_{e}(\theta_{e,t}-z_{e,t}),$ $\forall e,t$ (18)
@@ V-A Ablation Study
Figure 3: Ablation study illustrating the advantages of considering uncertainty, scouts for the reduction of uncertainty, and certainty decay for evolving conditions. Each plot shows the final paths of the robot units through the graph. Each unit is composed of a carrier robot and scout. The first case is the baseline from [2] and each subsequent case adds one of the components from the algorithm proposed in this paper to show the resulting effect to the final team routes through the graph. In (c) and (d), the darkness of the edges reflect the number of time steps the scouts explored that edge.
@@ VI CONCLUSION
In this paper, we explored multi-robot planning on dynamic topological graphs using mixed-integer programming for two challenging cases: heterogeneous teams and planning under uncertainty. We use the uncertainty in our problem as a motivation for a heterogeneous team and introduce scout robots that can investigate the environment to reduce the uncertainty of future team actions. Our approach results in a MILP problem that can be solved rapidly with a receding horizon in real-world scenarios. We tested this approach in example scenarios and demonstrated the ability to successfully generate risk-aware plans for multi-robot teams.